### v1

# 1D Spontaneous Symmetry Breaking in thermal equilibrium via Non-Hermitian Construction

Abstract
Spontaneous symmetry breaking generally circumvents one-dimensional systems with local interactions in thermal equilibrium. Here, we analyze a category of one-dimensional Hermitian models via local non-Hermitian constructions. Notably, spontaneous symmetry breaking and long-range order may emerge at finite temperatures in such systems under periodic boundary conditions, in sharp contrast to Hermitian constructions. We demonstrate clear numerical evidence, such as order parameters and specific heat, supporting phase diagrams with robust ordered phases. Non-Hermitian physics plays a vital role in prohibiting domain-wall proliferation and promoting spontaneous symmetry breaking. The fermions exhibit an exotic topological nature in their path-integral windings, which uphold nonzero integers - commonly a non-Hermitian signature - in the ordered phases, thus offering a novel and spontaneous origin for both symmetry breaking and non-Hermiticity.
Introduction— Spontaneous symmetry breaking (SSB) is a central concept in particle and condensed matter physics [1, 2, 3, 4, 5, 6, 7, 8], where the low-energy physics displays lower symmetry than the parent Hamiltonian, e.g., a ferromagnet, despite the parent Hamiltonian being spin symmetric, or a superconductor with broken ‘gauge’ symmetry, despite the parent Hamiltonian being manifestly gauge invariant. The Mermin-Wigner theorem has pointed out that due to proliferating defects and domain walls, SSB and long-range order do not materialize in one-dimensional (1D) systems with local interactions in thermal equilibrium [1, 2, 3, 8, 9, 10, 11, 12]. On the other hand, exceptions exist in classical systems far from thermal equilibrium, as SSB in 1D dynamical, diffusive and driven systems [13, 14, 15, 16, 17].
There has been much recent progress on non-Hermitian quantum systems, such as non-Hermitian topological phenomena [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31], the non-Hermitian skin effect (NHSE) [30, 32, 33, 34, 31], and nontrivial path-integral winding [35, 36]. While quantum mechanics dictates Hermitian Hamiltonians, the non-Hermitian models are increasingly proving their relevance in the fields of condensed matter physics [37, 38], finding their origins in open or dissipative systems [39, 40, 41, 42, 43, 44, 45, 46], optical [47, 48, 49, 50, 51, 52], acoustic [53, 54, 55, 56, 57, 58, 59, 60], electric [61, 62, 63, 64, 65, 66], mechanical [67, 68, 69, 70, 71, 72], and solid-state systems [73]. Recently, non-Hermitian physics of non-reciprocal classical fields has laid the foundations for understanding SSB in dynamical and non-equilibrium systems [17].
Here, we re-visit SSB in 1D quantum systems from the perspective of non-Hermitian constructions. Without loss of generality, we study 1D quantum systems where non-Hermitian fermions are locally coupled to a discrete field. The overall quantum systems are fully Hermitian and efficient for numerical and sign-problem-free Monte Carlo (MC) calculations. Interestingly, despite their 1D nature and local interactions, such systems exhibit clear signatures of long-range order following SSB, such as order parameters and specific heat behaviors, as the temperature is lowered. Therefore, for the first time, we deliver a 1D SSB scenario, which, unlike previous examples, is under thermal equilibrium and beyond classical dynamics.
The reasons that our constructions enable 1D SSB with local interactions and thermal equilibrium are as follows. From a energetic perspective, non-Hermitian systems exhibit unique non-locality - the NHSE shows that boundaries or domain walls impact their physics globally [30, 32, 33, 34, 31], and non-Hermitian delocalization yields long-range correlations and entanglements [35, 36] - therefore, non-Hermitian constructions may effectively introduce attractive strings between domain walls that contain their proliferation. From a fermionic perspective, the phase transition accompanies a characteristic and essentially topological shift of fermion path integrals from conventional ($w=0$) to nontrivial (simultaneous $w=\pm 2,\pm 4,\cdots$) path-integral winding-number sectors across the periodic boundary conditions (PBCs). Results from numerical calculations and mean-field theories also portray consistent physics. Although overall Hermitian, various non-Hermitian properties, e.g., such as nonzero winding number and nontrivial Greens functions [74], emerge in the ordered phase, exhibiting what we dub spontaneous ‘Hermiticity’ breaking - a novel route toward non-Hermitian physics.
Models— Without loss of generality, we consider the following 1D Hamiltonian under PBCs:
$\hat{H}$ $=$ $t\sum_{i}(c^{\dagger}_{i+1}c_{i}+c^{\dagger}_{i}c_{i+1})+t^{\prime}\sum_{i}(c^{\dagger}_{i+2}c_{i}+c^{\dagger}_{i}c_{i+2})$ (1)
$+$ $U\sum_{i}X_{i}(c^{\dagger}_{i+1}c_{i}-c^{\dagger}_{i}c_{i+1})-J\sum_{i}X_{i}X_{i+1},$
where $c_{i}$ is a fermion annihilation operator at site $x$, and $X_{i}=\pm 1$ takes discrete values on the bond between sites $i$ and $i+1$. $t=1$ and $t^{\prime}$ are the nearest-neighbor and next-nearest-neighbor hopping amplitudes of the fermions, respectively, $J$ is an Ising interaction between the nearest-neighbor $X_{i}$, and $U$ describes a local coupling between the fermions and $X_{i}$; see Fig. 1 for illustration.
We consider the property of $\hat{H}$ in thermal equilibrium at temperature $\beta=1/k_{B}T$, represented by its density operator $\hat{\rho}\propto e^{-\beta\hat{H}}$ and partition function:
$Z=\sum_{X}e^{-\beta E_{J}(X)}\cdot\mbox{Tr}_{c}[e^{-\beta\hat{H}(X)}]=\sum_{X}e^{-\beta E_{J}(X)}\prod_{n}[{1+e^{-\beta\varepsilon_{n}(X)}}],$ (2)
where $E_{J}(X)=-J\sum_{i}X_{i}X_{i+1}$, and $\varepsilon_{n}(X)$ is the spectrum of non-Hermitian fermion model $\hat{H}(X)$ following a specific ${X}$ configuration. For physical quantities’ expectation values, we average over $X$ with the probability:
$p({X})=\frac{1}{Z}e^{-\beta E_{J}(X)}\prod_{n}[{1+e^{-\beta\varepsilon_{n}(X)}}],$ (3)
which, given $\varepsilon_{n}(X)$ appears in complex-conjugate pairs, remains positive-definite and allows sign problem-free MC sampling. Further, through a similarity transformation, $\hat{H}(X)$ for $t^{\prime}=0$ maps to a translation-invariant Hatano-Nelson (HN) model, which depends only on the number $n_{\pm}$ of $X_{i}=\pm 1$ bonds; consequently, the fermion spectrum and weights are highly degenerate and the summation in Eq. 2 can be carried out numerically and exactly. Alternatively, we can analyze the model qualitatively via the self-consistently mean-field theory [74]:
$\hat{H}_{c}$ $=$ $\sum_{k}\left[2t\cos(k)+2t^{\prime}\cos(2k)+2iU\langle X\rangle_{\beta}\sin(k)\right]c^{\dagger}_{k}c_{k},$
$\hat{H}_{X}$ $=$ $X\left[\frac{2iU}{L}\sum_{k}\langle c^{\dagger}_{k}c_{k}\rangle_{\beta}\sin(k)-2J\langle X\rangle_{\beta}\right],$ (4)
where we have neglected the fluctuations of $X$ away from $\langle X\rangle_{\beta}$, and the $\langle\cdots\rangle_{\beta}$ are thermal expectation values. For further details of methods, see the Supplemental Material [74].
Figure 1: The model in Eq. 1 consists of fermions on the sites (black dots) and discrete $X_{i}=\pm 1$ (purple and cyan, respectively) on the bonds, which we sample or sum over for thermal equilibrium. The fermions’ nearest-neighbor hopping is non-reciprocal: $t\pm U$ to the right and $t\mp U$ to the left for $X_{i}=\pm 1$. More generically, we include an Ising interaction $J$ between nearest-neighbor $X_{i}$ and next-nearest-neighbor hopping $t^{\prime}$ of the fermions, which are not shown in the schematics.
Despite non-Hermitian $\hat{H}(X)$, the overall system remains Hermitian as we sum over $X$ configurations:
$\hat{\rho}=\frac{1}{Z}\sum_{X}\hat{\rho}(X)=\frac{1}{Z}\sum_{-X}\hat{\rho}(-X)=\frac{1}{Z}\sum_{X}\hat{\rho}^{\dagger}(X)=\hat{\rho}^{\dagger},$ (5)
where we have used $\hat{H}(-X)=\hat{H}^{\dagger}(X)$ and thus $\hat{\rho}(-X)=\hat{\rho}^{\dagger}(X)$ for the fermions given the $X$ configuration. The model is also invariant under translation as well as $X\rightarrow-X$ and $i\rightarrow-i$.
Figure 2: The order parameter $\langle X\rangle$ and specific heat $C_{V}$ indicate the presence of SSB and long-range order in the 1D models in Eq. 1 below a finite critical temperature $\beta>\beta_{C}$. The vertical dashed lines labels $\beta_{C}$ we pinpoint for $t^{\prime}=0.0$ through finite-size analysis (inset). Here, $U=0.4$, $J=0.0$, $t^{\prime}=0.0,0.5$, and $L=70$. We provide consistent MC and exact results for $t^{\prime}=0$, whereas only MC results are available at $t^{\prime}=0.5$.
Long-range order and energetics— We numerically evaluate physical properties such as the expectation value $\langle X\rangle$ and the specific heat $C_{V}$ over various system sizes and observe clear signatures of ordered phases at sufficiently low temperatures ($\beta>\beta_{C}$); see examples in Fig. 2. Here, $\langle X\rangle$ serves as an Ising-type order parameter and its long-range correlation emerges in the ordered phases [74]. The finite-size scaling of the specific heat $C_{V}$ accurately locates the transitions, which are essentially first order, as the preferential $\langle X\rangle$ the minimum of free energy $F(\langle X\rangle)$ jumps from one value to another without continuity [74]. We summarize the results in the phase diagrams in Fig. 3, where the ordered phases hold finite regions leaning towards larger $U$ and lower $T$, indicating their sufficient robustness towards variations of parameters. We note that aside from the first and ordering transition, there exists a series of additional transitions, whose physics we will explain later. On the contrary, the entire parameter spaces consist of only a disordered phase under open boundary conditions (OBCs) or without non-Hermitian construction, e.g., $U=0$ or $U=0.4i$ [74].
Figure 3: The phase diagrams in $U$ and $T$ contain finite regions of ordered phases at low $T$ and large $U$. Here, $t^{\prime}=0.0$ and $J=0,0.05$. The horizontal gray dashed line follows the model parameters in Fig. 2, which we use to determine $T_{C}$ ($\beta_{C}$). The solid curves are from the mean-field theory in Eq. 4 and Ref. [74].
Such outcomes are counter-intuitive, as SSB is generally absent in 1D with local interactions and thermalization, and the disordered phase should prevail at any finite temperature. The underlying reason is that the energy cost $E$ of an isolated domain wall is approximately constant  [^1] The constant energy cost results from the discrete symmetry, such as $X_{i}=\pm 1$ here, while the continuous symmetry introduces gapless Goldstone modes, making the long-range order even more prone towards destruction. and no match for its entropy $S$, which diverges with the system size $L\rightarrow\infty$ in the thermodynamic limit, in dominating the free energy $F=E-TS$ at arbitrarily low temperature $T$ [3]. Therefore, the domain walls proliferate, overthrow the long-range order, and forbid an SSB. Unlike previous 1D-SSB examples [13, 14, 16, 15] that circumvent the restrictions by staying far away from equilibrium, we make 1D SSB feasible via non-Hermitian construction’s unique non-local and energetic properties.
Let us consider the $t^{\prime}=0$ case in Eq. 1 as an example: given a specific $X$ configuration, the ground-state energy $E(X)$ of $\hat{H}(X)$ depends only on $n_{\pm}$; the minimum $E(X)$ corresponds to the uniform configuration $X_{0}$ with all $X_{i}=1$ ($X_{i}=-1$). Importantly, unlike the Ising interaction $E_{J}(X)$ that contributes a constant energy cost of $2J$ for each domain wall, $E(X)$, through non-Hermitian construction, is no longer locally dependent on the (number of) domain walls in $X$, but on the separations between them as well. For instance, $E(X)$ increases linearly with two domain walls $r=\alpha L$ apart, while $E(X)$ saturates to a constant with nearby domain walls in the thermodynamic limit, as we illustrate in Fig. 4. The trend remains at finite temperature. Such non-local energetic dependence applies to general scenarios beyond similarity transformations, e.g., $t^{\prime}\neq 0$, but are absent in similar yet Hermitian constructions or under OBCs [74].
Figure 4: The ground-state energy $E(X)$ increases linearly with two domain walls $r=\alpha L$ apart and saturates to a constant with nearby domain walls in the thermodynamic limit. Note the differences between the vertical scales. Here, $t^{\prime}=0,0.5$, $U=0.4$, and $X_{0}$ is uniformly ordered without domain walls. The insets illustrate the schematics of the corresponding $X$ configurations with two domain walls strung together by $E(X)$ and the constant energy cost $2J$ from $E_{J}$.
Such emergent non-local behaviors are unique to non-Hermitian physics and differ qualitatively from conventional Hermitian constructions, whose energy cost is only generated at or around the boundaries or domain walls  [^2] The NHSE is an example of how boundaries and domain walls impose global impacts; the non-Hermitian delocalization also introduces non-locality under and only under PBCs. . They are the key to the diminishing deterrence of SSB in 1D systems via non-Hermitian constructions: the domain walls are no longer free but strung together attractively proportional to their separations, prohibiting their global proliferation; despite recurrent fluctuations of domain-wall pairs in thermal equilibrium, they introduce merely local patches of opposing orders and pose no threat to the long-range order. We note that non-Hermitian constructions bear much potential and generalizability in modeling correlated physics beyond quasiparticles.
Topological aspects of fermion path integrals— We look into the fermion path integral in 1+1-dimensional space-(imaginary-)time underlying such SSB and ordered phases, once the $X$ field in Eq. 1 is integrated out. The consequences look trivial at first glimpse, as the periodicity in imaginary time $\beta$ mandates the fermions always to traverse an equal number of strengthened and attenuated bonds, and numerical results indeed eliminate the existence of nontrivial physics such as SSB or order phase under OBCs despite identical non-Hermitian construction [74]. Importantly, however, the fermion trajectories may circle the systems under PBCs, a topological nature labeled by an integer path-integral winding number $w$. The possibility of nonzero $w$ has been long overlooked, as the most relevant path integrals predominantly concentrate in the $w=0$ sector in conventional Hermitian models [77]. More recently, $w\neq 0$ sectors have begun to play vital roles, but only in non-Hermitian models. Nevertheless, as Hermitian conjugate maps $w$ to $-w$, having path integral simultaneously concentrated in the ±$\pm w$ sectors is equally valid in Hermitian quantum systems. Indeed, our phase diagram is characterized by such a topological nature: the SSB is signaled by the transition from the $w=0$ sector to the $w\neq 0$ sectors.
Figure 5: The phases and phase transitions exhibit a topological nature from the fermion perspective: its fermion path integral’s winding number $w=\operatorname{Im}(\langle v\rangle)\beta/L$ evolve from the conventional $w=0$ plateau in the disordered phase to nontrivial plateaus of $w=\pm 2,\pm 4,\cdots$ in the ordered phase. Here, $U=0.4$, $J=0.0$, and $t^{\prime}=0.0$. The vertical dashed line labels the corresponding transitions in Fig. 2. Inset: illustrations of trivial $w=0$ and nontrivial $w=2$ winding numbers in the fermion path integral under PBCs. Lower left: the probability distribution of $v$ shifts from peaking at 0 to finite-imaginary-value concentrations. Note the symmetry upon $v\rightarrow-v$, as Hermiticity requires. Lower right: long-range correlation of $v$ emerges in the ordered phases. The respective temperatures in the lower panels correspond to the colored arrows in the upper panel.
[…]
The SSB also leaves nontrivial marks on the fermions’ spectral functions: while singularities mainly concentrate on the real energy axis in the disordered phase, they emerge in imaginary regions on the complex energy plane in the order phases [74]. Such complex spectral properties, as well as the finite $\operatorname{Im}(\langle v\rangle)\neq 0$ in one of the SSB sectors, commonly relate to non-Hermitian physics - namely, such ordered systems have a spontaneously broken “Hermiticity”  [^4] The concept of spontaneous ‘Hermiticity’ breaking may sound unphysical at first glimpse. However, an analogy can be drawn with the broken ‘gauge’ symmetry in a superconducting phase - both Hermiticity and gauge invariance are fundamental requirements of a quantum theory and broken only at a mean-field level within a particular order-parameter sector, equivalent to other sectors via transformations, i.e., gauge transformations or Hermitian conjugate. . Such non-Hermitian features are also evident from a mean-field perspective, e.g., Eq. 4 with a finite $\langle X\rangle_{\beta}$.
Conclusions and discussion— Previous studies have found that Hermitian physics may emerge from non-Hermitian models or become influenced by non-Hermitian fixed points [80]. Here, we have shown that non-Hermitian construction may also impact the resulting Hermitian systems nontrivially, validating SSB even in one dimension, local interaction, and thermal equilibrium. On the one hand, such anomalies are due to the implicit non-locality induced by non-Hermitian physics, which correlates domain walls and upholds the long-range order; on the other hand, the phases and phase transitions are essentially topological from a quantum perspective, characterized by global integers of the fermion path-integral windings. Our numerical calculations have also provided accuracy and consistent results on the nature of such phases and SSB transitions. Thus, such non-Hermitian constructions carry novel prospects for SSB and non-Hermitian physics.
Non-Hermitian constructions are generalizable beyond our model examples to scenarios consisting of fermion or boson fields instead of $X$. The coupling similar to non-Hermitian $\hat{H}(X)$ will help introduce nontrivial, non-local quasiparticle energetics and thus establish exotic effective theory beyond free quasiparticles. Also, we obtain interacting, purely fermionic models by integrating out the $X_{i}=\pm 1$ degrees of freedom, which category we will discuss further in a separate future work. Interestingly, we have also demonstrated lively examples of a novel type of boundary conditions with a specific winding number $w$, where the $w=0$ case is simply OBCs  [^5] Such boundary conditions are dual to the flux phases $e^{i\phi}$ across PBCs. . Such constructions and boundary conditions offer a unique and non-perturbative variational dimension, which may significantly impact the universal physics of condensed matter systems [74].
Acknowledgment: We acknowledge insightful discussions with Yongxu Fu and generous support from the National Natural Science Foundation of China (Grants No.92270102 & No.12174008) and the National Key R&D Program of China (Grant No.2022YFA1403700).